# Reinforcement Learning Enhancing Entanglement for Two-Photon-Driven Rabi Model

### v3

@@ IV.4 Entanglement under control
We check the dynamics of the entanglement witness $E^{T}$ under the control versus different system parameters and initial states. The time interval [0, 1] is partitioned into 30 segments of control square pulses with amplitudes picked in $\{0,~{}0.5\Omega_{max}/\delta_{c},~{}\Omega_{max}/\delta_{c}\}$ determined by the RL agent. The RL agent designs the arrangement of the pulses under the objective of increasing the entanglement. Each round of the training consists of 100 epochs. As an example, one round of the control process to find the sequence of square pulses is shown in Fig. 5. As shown in Fig. 5($a$), along with the training, the RL agent can find numerous sequences of square pulses all with increased entanglement. The corresponding control sequences are shown in Fig. 5($b$). This control method can be regarded as a kind of combination of bang-bang control PRA582733 ; PRA86022321 and reinforcement learning. From a statistical view, it can be intuitively observed that as learning progresses, the probability of $\Omega_{max}/\delta_{c}=0$ decreases in the distribution of the control pulses. These solutions corroborate the efficacy of the control pulses designed by the RL agent to enhance the generation of entanglement.
@@
Figure 6: The average evolutions of 20 entanglement criterion $E^{T}$ with error zones for different two-photon-driven amplitudes and initial states. $(a_{1})$ $\Omega_{max}=0.3\delta_{c}$ and $\rho(0)=\rho_{0}$, the ground state of the Hamiltonian (1). $(a_{2})$ $\Omega_{max}=0.3/\delta_{c}$ and $\rho(0)$= $|\alpha=1,\downarrow\rangle\langle\alpha=1,\downarrow|$. $(b_{1})$ $\Omega_{max}=0.5\delta_{c}$ and $\rho(0)=\rho_{0}$ is the ground state of the Hamiltonian (1). $(b_{2})$ $\Omega_{max}=0.5\delta_{c}$ and $\rho(0)$= $|\alpha=1,\downarrow\rangle\langle\alpha=1,\downarrow|$. The other parameters are same to those in Fig. 5.

### v4

@@ IV.4 Entanglement under control
We conjecture the amplitude of the control field may have effect on the enhancing effect. For the initial states $\rho(0)=\rho_{0}$, namely, the ground state of the Hamiltonian (1) and $\rho(0)$ = $|\alpha=1,\downarrow\rangle\langle\alpha=1,\downarrow|$, we check the influence of the maximum amplitude $\Omega_{max}/\delta_{c}$ on the optimized entanglement-enhancing case in Fig. 5. This provides a reference for selecting the value of $\Omega_{max}/\delta_{c}$. In the following research, we mainly focus on $\Omega_{max}/\delta_{c}$, which is a moderate choice with respect to the system parameters.
We check the dynamics of the entanglement witness $E^{T}$ under the control versus different system parameters and initial states. The time interval [0, 1] is partitioned into 30 segments of control square pulses with amplitudes picked in $\{0,~{}0.5\Omega_{max}/\delta_{c},~{}\Omega_{max}/\delta_{c}\}$ determined by the RL agent. The RL agent designs the arrangement of the pulses under the objective of increasing the entanglement. Each round of the training consists of 100 epochs. As an example, one round of the control process to find the sequence of square pulses is shown in Fig. 6. As shown in Fig. 6($a$), along with the training, the RL agent can find numerous sequences of square pulses all with increased entanglement. The corresponding control sequences are shown in Fig. 6($b$). This control method can be regarded as a kind of combination of bang-bang control PRA582733 ; PRA86022321 and reinforcement learning. From a statistical view, it can be intuitively observed that as learning progresses, the probability of $\Omega_{max}/\delta_{c}=0$ decreases in the distribution of the control pulses. Yet, the negative effect of dissipation on control performance can be observed in Fig. 6($c$) and ($d$). With the same other parameters, the stronger the dissipation, the worse the effect of RL enhanced entanglement. These solutions corroborate the efficacy of the control pulses designed by the RL agent to enhance the generation of entanglement.
Figure 6: (a) The dynamics of the entanglement indicated by $E^{T}$ versus the training epoches. (b) The corresponding square control pulses $\Omega(t)$ ($\Omega_{max}/\delta_{c}$ = 0.3) from the top view, with values shown in the color bar. $\rho(0)=\rho_{0}$ is the ground state of the Hamiltonian (1), $\delta_{q}/\delta_{c}$ = 1, $g/\delta_{c}$ = 0.3, $K/\delta_{c}$ = 0, and dissipation rates $\gamma_{\hat{a}}/\delta_{c}$ = $\gamma_{\hat{\sigma}_{-}}/\delta_{c}$ = 0.01. (c) and (d) Evolution of entanglement witness $E^{T}$ for different dissipations as shown in the legend: $\gamma_{c}/\delta_{c}$ = $\gamma_{q}/\delta_{c}$ = 0.01, 0.03, 0.05, 0.07, 0.09, for $\rho(0)=\rho_{0}$, the ground state of the Hamiltonian (1) and $\rho(0)$ = $|\alpha=1,\downarrow\rangle\langle\alpha=1,\downarrow|$, respectively.
@@
Figure 7: The average evolutions of 20 entanglement criterion $E^{T}$ with error zones for different two-photon-driven amplitudes and initial states. $(a_{1})$ $\Omega_{max}=0.3\delta_{c}$ and $\rho(0)=\rho_{0}$, the ground state of the Hamiltonian (1). $(a_{2})$ $\Omega_{max}=0.3/\delta_{c}$ and $\rho(0)$ = $|\alpha=1,\downarrow\rangle\langle\alpha=1,\downarrow|$. $(b_{1})$ $\Omega_{max}=0.5\delta_{c}$ and $\rho(0)=\rho_{0}$ is the ground state of the Hamiltonian (1). $(b_{2})$ $\Omega_{max}=0.5\delta_{c}$ and $\rho(0)$ = $|\alpha=1,\downarrow\rangle\langle\alpha=1,\downarrow|$. The other parameters are the same to those in Fig. 6.